OpenStudy (anonymous):

lim as x approaches 0 of (x-1)/(x^3-2x^2)

OpenStudy (raden):

directly subtittute x = 0 into the function of (x-1)/(x^3-2x^2) what do you get ?

OpenStudy (anonymous):

undefined!

OpenStudy (anonymous):

but the answer is infinity

OpenStudy (raden):

(0-1)/(0^3-2(0)^2) = -1/(0-0) = -1/0 = -infinity

OpenStudy (raden):

yes, it is infinity. just miss the negative sign above

OpenStudy (anonymous):

but can't you not divide by 0? so wouldn't i not be allowed to do that?

OpenStudy (raden):

a number divide by 0 would be infinity, but 0/0 its undefined

OpenStudy (anonymous):

okay thank you!!!

OpenStudy (raden):

wlcm